# Infinite-Dimensional Closed-Loop Inverse Kinematics for Soft Robots via Neural Operators

### v1

@@ Abstract
While kinematic inversion is a purely geometric problem for fully actuated rigid robots, it becomes extremely challenging for underactuated soft robots with infinitely many degrees of freedom. Closed-loop inverse kinematics (CLIK) schemes address this by introducing end-to-end mappings from actuation to task space for the controller to operate on, but typically assume finite dimensions of the underlying virtual configuration space. In this work, we extend CLIK to the infinite-dimensional domain to reason about the entire soft robot shape while solving tasks. We do this by composing an actuation-to-shape map with a shape-to-task map, deriving the differential end-to-end kinematics via an infinite-dimensional chain rule, and thereby obtaining a Jacobian-based CLIK algorithm. Since the actuation-to-shape mapping is rarely available in closed form, we propose to learn it from simulation data using neural operator networks, which are differentiable. We first present an analytical study on a constant-curvature segment, and then apply the neural version of the algorithm to a three-fiber soft robotic arm whose underlying model relies on morphoelasticity and active filament theory. This opens new possibilities for differentiable control of soft robots by exploiting full-body shape information in a continuous, infinite‑dimensional framework.
@@ I Introduction
Kinematic inversion refers to finding a feasible robot configuration such that the robot assumes a desired positioning of one or more of its parts in space [24]. For a rigid manipulator, this usually translates to computing joint positions that result in a desired end-effector pose, given the geometric link parameters. This formulation, however, assumes full actuation of rigid robots. When the robot itself becomes soft and can bend and twist continuously rather than only at discrete joints, its nature-inspired flexibility opens new possibilities in navigating complex environments and interacting safely with delicate objects, yet this same property makes control in general challenging and even the usually straightforward task of kinematic inversion difficult.
Related Work. For fully actuated systems, where the degrees of actuation equal the degrees of freedom, kinematic inversion is a purely geometric problem, in which desired joint configurations can be selected freely from the entire configuration space. Jacobian-based closed-loop inverse kinematics (CLIK) schemes exploit this by using feedback in task-space to map back to required joint velocities via the Jacobian inverse, avoiding the need to solve the full nonlinear inverse kinematics problem at every step [32, 15, 21, 9]. However, in under-actuated systems with fewer degrees of actuation than degrees of freedom, such as soft robots, not all configurations are attainable through control action [14]. Static kinematic constraints must be enforced explicitly, typically via a numerical optimization, which is then embedded, for example, in a model predictive control fashion [37, 12, 34, 36]. A compact, closed-loop form similar to the fully actuated case only becomes possible once we introduce an end-to-end kinematic mapping from the actuator inputs to the task-space output [38]: This mapping plays the role of a virtual joint-space model, like a controllable subspace, on which standard CLIK algorithms can then operate [17, 6, 13, 2, 31]. This so-called actuator-to-task closed-loop inverse kinematics approach was formalized for underactuated mechanical systems in [11], however, still assuming finite dimensions for the underlying virtual joint space. i.e., relying on the common soft robot approximation that the shape is linearly parametrizable with a set of basis functions. But what if we want to reason on the entire (infinite-dimensional) soft robot shape while solving tasks?
@@
Contribution. In this work, we extend CLIK to systems with infinite-dimensional shape spaces, while keeping the controller finite-dimensional through end-to-end mappings from finite actuators to finite task dimensions. To this end, we define two mappings: (i) the mapping from actuators to shape, that is, the “internal” model of the (soft) robot, (ii) the forward kinematics from an infinite-dimensional shape space to a task, whose composition yields the end-to-end forward kinematics from actuation to task. Combining the gradients of both mappings via an infinite-dimensional chain rule yields the end-to-end forward differential kinematics, and the standard Jacobian-based CLIK algorithm applies. It quickly becomes obvious that the key challenge here is that soft robots’ actuator-to-shape mappings are highly complex and often lack closed-form expressions or explicit Jacobians, so that the proposed CLIK algorithm is only practical with a learned, differentiable model of this map. To adress this, we propose to learn the mapping from simulation data using a neural operator network, which, unlike traditional neural networks, learns mappings between infinite-dimensional function spaces without fixed discretization, while providing well-defined gradients with respect to the finite-dimensional inputs [1, 27]. Neural operators are a recent trend in machine learning that have found wide interest in control and (soft) robotics already, for example as surrogate models for complex dynamics and soft materials [5, 16, 26, 28, 7], or to make complex or computationally expensive control schemes either real-time feasible or differentiable [3, 22, 23, 10, 18]. For us, they enable reasoning about the full infinite-dimensional soft robot shape during tasks.
Organization. First, we review finite-dimensional CLIK for underactuated mechanical systems in Section II, and then introduce its extension to infinite-dimensional shapes in Section III. Last but not least, Section IV presents the neural version of the algorithm.
II Preliminaries: (Finite-Dimensional) Closed-Loop Inverse Kinematics
Before introducing our extension to infinite-dimensional closed-loop inverse kinematics, we will briefly review the finite-dimensional case and unify notations. We refer to [33] and [11] for extended clarifications.
Kinematic inversion. Intuitively, forward kinematics answer “If I set the robot joints to these angles, where does the tip end up?”, while kinematic inversion solves the reverse: “Which joint angles will get my tip to the desired position?” Formally, let $q\in\mathbb{R}^{n}$ be the configuration space of a mechanical system with $n$ degrees of freedom. Further, let $x\in\mathbb{R}^{p}$ be a $p$-dimensional task coordinate, commonly the end-effector position and orientation, such that the robot fulfilling the tasks is expressed through the forward kinematics
@@ II Preliminaries: (Finite-Dimensional) Closed-Loop Inverse Kinematics
Closed-loop inverse kinematics (CLIK). This control scheme looks for solutions of the kinematic inversion in the form of dynamical system
$\dot{q}$ $=c_{\text{IK}}(q,\bar{x})$
$\text{s.t.}\lim_{t\to\infty}\Phi_{\text{fin}}(q(t))=\bar{x}$ (2)
so that its equilibrium $\bar{q}$ satisfies the kinematic inversion $\bar{x}=\Phi_{\text{fin}}(\bar{q})$. A standard choice for this dynamical system that ensures exponential convergence under mild assumptions is
@@
with positive definite gain matrix $K$ and associated Jacobian $J_{\Phi_{\text{fin}}}$, which is square and invertible for the fully actuated case with $p=n$. It follows from linearizing the task error $\tilde{x}=\bar{x}-\Phi_{\text{fin}}(q)$ and requiring the closed-loop dynamics $\dot{\tilde{x}}=-K\tilde{x}$ in task space. Inverting through the Jacobian yields  (3), which exhibits exponential convergence provided the Jacobian remains invertible.
CLIK for underactuated systems. For underactuated systems with $m<n$, we cannot control the $n$-dimensional configuration space, but we can extend the above to finding a control $c_{\text{IK}}$ that also respects static constraints, i.e.,
@@
where $\mathcal{E}$ is the attainable equilibria set containing all configurations that can be reached as a steady state [11]. To realize this, we have to introduce an end-to-end mapping from actuators to task output that creates a virtual configuration subspace controllable by actuators on which CLIK can operate. Therefore, consider the finite-dimensional mapping
@@
from actuation coordinates $q_{\text{a}}$ with $m$ independent degrees of actuation to configuration coordinates $q$, such that the composed forward pseudo-kinematics that maps actuators to task variables are
@@
Note that, here, we focus on a square inversion problem where the actuation space and task space have equal dimension ($m=p$), making the actuation-to-task mapping non-redundant. The corresponding square Jacobian is $J_{(\Phi\circ\Lambda)_{\text{fin}}}(q_{\text{a}})$. Then, the standard gradient-based CLIK algorithm is
@@
where a positive definite gain matrix $K\in\mathbb{R}^{m\times m}$ ensures exponential convergence analogous to the fully actuated case.
@@ III Infinite-Dimensional Closed-Loop Inverse Kinematics
For the infinite-dimensional analogue of (7), we introduce a general (soft) robot model via the actuation-to-shape mapping
@@ III-A Differential Kinematics
We observe that the resulting error dynamics $\dot{\bar{x}}-\dot{x}=-K(\bar{x}-x)$ is linear and converges exponentially fast to zero with the rated defined by $K$. Consequently, the steady state of (23) fulfills $\lim_{t\to\infty}\Phi(q(t))=\bar{x}$. ∎
@@ III-B Representative Tasks and Resulting Jacobians
meaning that the closest centerline coordinate $s_{*}$ to the target is determined through optimization at each step.
@@
where, the closest centerline coordinate $s_{*}$ directly results from the distance optimization. Assuming uniqueness and nondegeneracy of the minimizer (Infinite-dimensional Danskin’s Theorem, cf. Theorem 4.13 in [4]), its directional derivative follows (15) with $L^{2}$-gradients $\nabla_{L^{2}}\Phi_{\text{dist}}^{\text{fixed}}(r)=(r(s_{*})-x_{0})\delta(s-s_{*})\in X$, and $\nabla_{L^{2}}\Phi_{\text{dist}}^{\text{opt}}(r)=(r(s_{*})-x_{0})\delta(s-\bar{s})\in X$, such that the row Jacobians of the end-to-end mapping $J_{\Phi\circ\Lambda}\in\mathbb{R}^{1\times m}$ have $m$ scalar entries
@@ IV-A Neural Operator Networks and Their Gradients
In practice, the branch network is trained on finitely many function samples, while at inference time the trunk network can be evaluated ar arbitrarily many, arbitrarily fine locations $s$. This decouples the discretization of the input function from the resolution of the output, which is a key advantage of neural operators over discretization-based neural networks, and makes it particularly powerful for infinite-dimensional CLIK where we want to reason over the continuous robot shape and might be interested in finer discretizations.
@@ IV-B Example: Three-fiber Soft Robotic Arm
To demonstrate our neural infinite-dimensional CLIK algorithm, we consider a soft robotic arm initially designed with three liquid crystal elastomer fibers, inspired by the elephant trunk musculature. It is depicted in Figure 5. By design, this structure is soft and slender, and can grow, shrink, or actively remodel its intrinsic shape while also deforming elastically – which is why it is modelled as an active, morphoelastic filament [29, 20]. This model has been previously validated in experiments [25, 19], successfully used for path-planning algorithms [35], and shall now serve as example for the neural operator implementation of infinite-dimensional CLIK. We briefly summarize this model here, but refer to [20, 19] for details.
Morphoelasticity. Consider a tubular body of length $L$, which is reduced to the centerline $\bm{r}:[0,1]\rightarrow\mathbb{R}^{3}$ with normalized spatial coordinate $s\in[0,1]$, and an orthonormal director frame $D=\{d_{1},d_{2},d_{3}\}$, where $d_{i}:[0,L]\rightarrow\mathbb{R}^{3}$, is attached to $r(s)$ for all $s$. The kinematics of the one-dimensional filament follow as
$r^{\prime}(s)$ $=\zeta d_{3},$ (42a)
$d_{i}^{\prime}(z)$ $=\zeta u\times d_{i},\quad i=\{1,2,3\},$ (42b)
with the axial extension of the filament $\zeta$ and the Darboux curvature vector $u=u_{1}d_{1}+u_{2}d_{2}+u_{3}d_{3}$ describing the evolution of the director basis $D$ along the filament. Without external load, we refer to these properties as intrinsic curvatures $u=\hat{u}$ and intrinsic extension $\zeta=\hat{\zeta}$. Under (gravitational) load, intrinsic and actual properties differ, and we further need to consider the quasi-static solution of the force and moment balance equations, namely
$\frac{\partial n}{\partial s}+\hat{\zeta}f=0,\quad\frac{\partial m}{\partial s}+\frac{\partial r}{\partial s}\times n+\hat{\zeta}l=0.$ (43)
Here, $n$ is the internal force, $\hat{\zeta}f$ is the external body force per unit length, $m$ is the internal momentum, and $\hat{\zeta}l$ is the external body couple per unit length.
Active filaments. To connect the filament model to the robot’s actuation, we treat the soft robot as an active filament whose material can locally shorten or extend along embedded fibers. The internal fiber architecture is described by a fiber direction field inside the filament body, which may be uniaxial or helical, and the corresponding intrinsic properties are obtained by integrating the fiber contributions over the cross section. While the proposed approach allows for general forms of activation $\gamma_{i}$, in this work we consider piecewise uniform activations and obtain a single scalar value $\gamma_{i}\in\mathbb{R}$ per fiber. We identify these activation scalars with the actuation coordinates and write $q_{\text{a},i}:=\gamma_{i}$, so that the actuation coordinate directly scales the intrinsic contraction of extension generated by each fiber. Negative activation $\gamma_{i}<0$ (or $q_{\text{a},i}<0$) corresponds to contraction, positive activation $\gamma_{i}>0$ (or $q_{\text{a},i}>0$) corresponds to extension.
Three-fiber soft robotic arm. The soft robot considered in this work has one longitudinal and two helical fibers inspired by the elephant trunk musculature (Figure 5) and is subject to gravitational load. Combining the above, its forward kinematics are governed by
@@
where the mapping from actuation $q_{\text{a}}\in\mathbb{R}^{3}$ to shape $r(s)$ is mediated by the intrinsic curvature $\hat{u}$ and extension $\hat{\zeta}$, both of which are functions of the fiber activations, hence, the actuation, i.e., $\hat{u}(q_{\text{a}})$ and $\hat{\zeta}(q_{\text{a}})$.
@@ IV-C Learning the Model with a Neural Operator
In our application, the architecture of the operator network is simplified compared to the DeepONet architecture [27], since the input is a single scalar-valued actuation function and the actuation coordinates $q_{\text{a}}\in\mathbb{R}^{m}$ are already finite-dimensional. Consequently, the branch net reduces to a standard feedforward network that takes $q_{\text{a}}$ as input, while the trunk net only encodes the spatial variable $s$. At the same time, we retain the key property that the operator is learned over actuation and space, so we can evaluate the learned actuation-to-shape mapping $\mathcal{G}_{\Lambda}$ at arbitrarily many points $s$ and differentiate with respect to $q_{\text{a}}$, which is exactly what the neural CLIK algorithm requires.
@@ IV-D Composed Jacobian and Neural CLIK
Figure 6 shows simulations of the neural version of the infinite-dimensional CLIK algorithm for the two distinct positioning tasks. The simulation is for $t\in[0,1]$ with a time step of $dt=0.001$. The weighting matrix is $K=8\cdot\text{diag}(1,1,1)$, i.e., all positioning coordinates are equally weighted, and the magnitude is a balance between convergence speed and overshooting. We use a spatial discretization with $n_{\text{s}}=100$ points in this example, but this discretization could be arbitrarily finer. The code will be made available on GitHub upon acceptance.

### v2

@@ Abstract
For fully actuated rigid robots, kinematic inversion is a purely geometric problem, efficiently solved by closed-loop inverse kinematics (CLIK) schemes that compute joint configurations to position the robot body in space. For underactuated soft robots, however, not all configurations are attainable through control action, making kinematic inversion extremely challenging. Extensions of CLIK address this by introducing end-to-end mappings from actuation to task space for the controller to operate on, but typically assume finite dimensions of the underlying virtual configuration space. In this work, we formulate CLIK in the infinite-dimensional domain to reason about the entire soft robot shape while solving tasks. We do this by composing an actuation-to-shape map with a shape-to-task map, deriving the differential end-to-end kinematics via an infinite-dimensional chain rule, and thereby obtaining a Jacobian-based CLIK algorithm. Since this actuation-to-shape mapping is rarely available in closed form, we propose to learn it using differentiable neural operator networks. We first present an analytical study on a constant-curvature segment, and then apply the neural version of the algorithm to a three-fiber soft robotic arm whose underlying model relies on morphoelasticity and active filament theory.
@@ I Introduction
Kinematic inversion refers to finding a feasible robot configuration such that the robot assumes a desired positioning of one or more of its parts in space. For a rigid manipulator, this translates to computing joint positions that result in a desired end-effector pose, given the geometric link parameters. This formulation, however, assumes full actuation of rigid robots. When a robot is soft, it can bend and twist continuously rather than only at discrete joints. This flexibility enables navigation in complex environments and safe interaction with delicate objects, but it also makes control, even basic kinematic inversion, challenging.
For fully actuated systems, Jacobian-based closed-loop inverse kinematics (CLIK) schemes exploit the geometric freedom to select joint configurations directly, by using feedback in task-space to map back to required joint velocities via the Jacobian inverse [11, 16, 7]. However, in under-actuated systems with fewer degrees of actuation than degrees of freedom, such as soft robots, not all configurations are attainable through control action. Static kinematic constraints must be enforced explicitly, typically via a numerical optimization, which is then embedded, for example, in a model predictive control fashion [25, 24]. A compact, closed-loop form similar to the fully actuated case only becomes possible once we introduce an end-to-end kinematic mapping from the actuator inputs to the task-space output [26]: This mapping plays the role of a virtual joint-space model, like a controllable subspace, on which standard CLIK algorithms can then operate [5, 10, 21]. This so-called actuator-to-task closed-loop inverse kinematics approach was formalized for underactuated mechanical systems in [9]. Yet, still assuming finite dimensions for the underlying virtual joint space. i.e., relying on the common soft robot approximation that the shape is linearly parametrizable with a set of basis functions.
@@
However, many soft-robot tasks require reasoning about the entire body shape, such as positioning any suitable point to a target or enforcing contact constraints along the body. In this work, we extend CLIK to systems with infinite-dimensional shape spaces, while keeping the controller finite-dimensional through end-to-end mappings from finite actuators to finite task dimensions. To this end, we define two mappings: (i) the mapping from actuators to shape, that is, the “internal” model of the (soft) robot, (ii) the forward kinematics from an infinite-dimensional shape space to a task, whose composition yields the end-to-end forward kinematics from actuation to task. Combining the gradient operators of both mappings via an infinite-dimensional chain rule yields the end-to-end differential forward kinematics, enabling the formulation of Jacobian-based CLIK. Since soft robots’ actuator-to-shape mappings often lack closed-form expressions or explicit Jacobians, this is only practical with a learned, differentiable model of this map. To address this, we propose to learn the mapping from simulation data using a neural operator network, which, unlike traditional neural networks, learns mappings between infinite-dimensional function spaces without fixed discretization, while providing well-defined gradients with respect to the finite-dimensional inputs [1, 18]. Neural operators are a recent trend in machine learning that have found wide interest in control and (soft) robotics already, for example as surrogate models for complex dynamics [4, 12, 19], or to make computationally expensive control schemes either real-time feasible or differentiable [2, 8, 13]. For us, they enable reasoning about the full infinite-dimensional soft robot shape during tasks.
II Preliminaries: Finite-Dimensional Closed-Loop Inverse Kinematics
We will briefly review the finite-dimensional case and unify notations, but refer to [22, 9] for details.
Kinematic inversion. Let $q\in\mathbb{R}^{n}$ be the configuration space of a mechanical system with $n$ degrees of freedom. Further, let $x\in\mathbb{R}^{p}$ be a $p$-dimensional task coordinate, commonly the end-effector position and orientation, such that the robot fulfilling the tasks is expressed through the forward kinematics
@@ II Preliminaries: Finite-Dimensional Closed-Loop Inverse Kinematics
Closed-loop inverse kinematics (CLIK). CLIK solves the kinematic inversion in the form of dynamical system
$\dot{q}=c_{\text{IK}}(q,\bar{x})\qquad\text{s.t.}\lim_{t\to\infty}\Phi_{\text{fin}}(q(t))=\bar{x}.$ (2)
A standard choice that ensures exponential convergence under mild assumptions is
@@
with positive definite gain matrix $K$ and Jacobian $J_{\Phi_{\text{fin}}}$, which is square and invertible for the fully actuated case.
CLIK for underactuated systems. For underactuated systems, we cannot control the $n$-dimensional configuration space, but we can extend the above to finding a control $c_{\text{IK}}$ that also respects static constraints, i.e.,
@@
where $\mathcal{E}$ is the attainable equilibria set. To realize this, we introduce an end-to-end mapping from actuators to task output that creates a virtual configuration subspace controllable by actuators on which CLIK can operate. Therefore, consider the finite-dimensional mapping
@@
from actuation coordinates $q_{\text{a}}$ with $m$ independent degrees of actuation to configuration coordinates $q$, such that the composed forward pseudo-kinematics are
@@
Considering a square inversion ($m=p$) with Jacobian $J_{(\Phi\circ\Lambda)_{\text{fin}}}(q_{\text{a}})$, the standard gradient-based CLIK algorithm is
@@
A positive definite gain matrix $K\in\mathbb{R}^{m\times m}$ ensures exponential convergence analogous to the fully actuated case.
@@ III Infinite-Dimensional Closed-Loop Inverse Kinematics
For the infinite-dimensional analogue of (7), we introduce a (soft) robot model via the actuation-to-shape mapping
@@ III-A Differential Kinematics
We observe that the resulting error dynamics $\dot{\bar{x}}-\dot{x}=-K(\bar{x}-x)$ is linear and converges exponentially fast to zero with the rate defined by $K$. Consequently, the steady state of (23) fulfills $\lim_{t\to\infty}\Phi(q(t))=\bar{x}$. ∎
@@ III-B Representative Tasks and Resulting Jacobians
meaning that the closest centerline coordinate $s_{*}$ to the target is computed via gradient descent in each step.
@@
where, the closest centerline coordinate $s_{*}$ is computed via gradient descent. Assuming uniqueness and nondegeneracy of the minimizer (Infinite-dimensional Danskin’s Theorem, cf. Theorem 4.13 in [3]), its directional derivative follows (15) with $L^{2}$-gradients $\nabla_{L^{2}}\Phi_{\text{dist}}^{\text{fixed}}(r)=(r(s_{*})-x_{0})\delta(s-s_{*})\in X$, and $\nabla_{L^{2}}\Phi_{\text{dist}}^{\text{opt}}(r)=(r(s_{*})-x_{0})\delta(s-\bar{s})\in X$, such that the row Jacobians of the end-to-end mapping $J_{\Phi\circ\Lambda}\in\mathbb{R}^{1\times m}$ have $m$ scalar entries
@@ IV-A Neural Operator Networks and Their Gradients
In practice, the branch network is trained on finitely many function samples, while at inference time the trunk network can be evaluated at arbitrarily many, arbitrarily fine locations $s$. This decouples the discretization of the input function from the resolution of the output, which is a key advantage of neural operators over discretization-based neural networks, and makes it particularly powerful for infinite-dimensional CLIK where we want to reason over the continuous robot shape and might be interested in finer discretizations.
@@ IV-B Example: Three-fiber Soft Robotic Arm
To demonstrate our neural infinite-dimensional CLIK algorithm, we consider a soft robotic arm initially designed with three contractable fibers, inspired by the elephant trunk musculature. It is depicted in Figure 5. By design, this structure is soft and slender, and can grow, shrink, or actively remodel its intrinsic shape while also deforming elastically – which is why it is modelled as an active, morphoelastic filament [15]. This model has been previously validated in experiments [17, 14], successfully used for path-planning algorithms [23], and shall now serve as example for the neural operator implementation of infinite-dimensional CLIK. Its forward kinematics are governed by
@@
with centerline $r:[0,1]\rightarrow\mathbb{R}^{3}$, normalized spatial coordinate $s\in[0,1]$, orthonormal director frame $D=\{d_{1},d_{2},d_{3}\}$, as well as extensions $\zeta$, curvatures $u$, internal force $n$, external force $f$, internal momentum $m$. The mapping from actuation $q_{\text{a}}\in\mathbb{R}^{3}$ to shape $r(s)$ is mediated by the intrinsic curvature $\hat{u}$ and extension $\hat{\zeta}$, both of which are functions of the fiber activations, hence, the actuation, i.e., $\hat{u}(q_{\text{a}})$ and $\hat{\zeta}(q_{\text{a}})$. Details are found in [15, 14].
@@ IV-C Learning the Model with a Neural Operator
In our application, the architecture of the operator network is simplified compared to DeepONet [18], since the actuation coordinates $q_{\text{a}}\in\mathbb{R}^{m}$ are finite-dimensional. Consequently, the branch net reduces to a standard feedforward network, while the trunk net only encodes the spatial variable $s$. We still retain the key property that the operator is learned over actuation and space, so we can evaluate the learned actuation-to-shape mapping $\mathcal{G}_{\Lambda}$ at arbitrarily many points $s$ and differentiate with respect to $q_{\text{a}}$.
@@ IV-D Composed Jacobian and Neural CLIK
Figure 6 shows simulations of the neural version of the infinite-dimensional CLIK algorithm for the two distinct positioning tasks. The simulation is for $t\in[0,1]$ with a time step of $dt=0.001$. The weighting matrix is $K=8\cdot\text{diag}(1,1,1)$, i.e., all positioning coordinates are equally weighted, and the magnitude is a balance between convergence speed and overshooting. We use a spatial discretization with $n_{\text{s}}=100$ points in this example, but this discretization could be arbitrarily finer. The code and dataset will be made available on GitHub upon acceptance.